# Linear Mode Connectivity in Differentiable Tree Ensembles

### v1

@@ Appendix B Dataset
Appendix C Additional Empirical Results
@@
Figures 9, 10, 11, 12, 13, 14, 15 and 16 show the interpolation curves of oblivious trees with $D=2$ and $M=256$ across various datasets and configurations. Significant improvements are particularly noticeable in AM, but improvements are also observed in WM. These characteristics are also apparent in the non-oblivious trees, as shown in Figures 17, 18, 19, 20, 21, 22, 23 and 24. Regarding split data training, the dataset for each of the two classes is initially complete (100%). It is then divided into splits of 80% and 20%, and 20% and 80%, respectively. Each model is trained using these splits. Figures 13, 15, 21, and 23 show the training accuracy evaluated using the full dataset (100% for each class).

### v2

@@ Appendix A Dataset
Appendix B Additional Empirical Results
@@
Figures 10, 11, 12, 13, 14, 15, 16 and 17 show the interpolation curves of oblivious trees with $D=2$ and $M=256$ across various datasets and configurations. Significant improvements are particularly noticeable in AM, but improvements are also observed in WM. These characteristics are also apparent in the non-oblivious trees, as shown in Figures 18, 19, 20, 21, 22, 23, 24 and 25. Regarding split data training, the dataset for each of the two classes is initially complete (100%). It is then divided into splits of 80% and 20%, and 20% and 80%, respectively. Each model is trained using these splits. Figures 14, 16, 22, and 24 show the training accuracy evaluated using the full dataset (100% for each class). In split data training, the performance reference of full data training is shown only for the performance on the test data. This is because, in split data training, even the training dataset used for evaluation includes portions that are not used for training each model, which differs from the conditions in full data training. In contrast, when evaluating performance on the test data, all of the test data have not been used equally for the training of each model, which allows for a fair comparison between the two approaches. Figures 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40 and 41 visualize the same to that of perfect binary trees for the decision lists.
Figures 42 and 43 show the interpolation curves for MNIST [35] with various tree architectures where $D=2$ and $M=256$. Although MNIST consists of 2-dimensional image data, it is input as a 1-dimensional vector.
@@
Figure 26: Interpolation curves of train accuracy for decision lists with AM.
Figure 27: Interpolation curves of test accuracy for decision lists with AM.
Figure 28: Interpolation curves of train accuracy for decision lists with WM.
Figure 29: Interpolation curves of test accuracy for decision lists with WM.
Figure 30: Interpolation curves of train accuracy for decision lists with AM by use of split dataset.
Figure 31: Interpolation curves of test accuracy for decision lists with AM by use of split dataset.
Figure 32: Interpolation curves of train accuracy for decision lists with WM by use of split dataset.
Figure 33: Interpolation curves of test accuracy for decision lists with WM by use of split dataset.
Figure 34: Interpolation curves of train accuracy for modified decision lists with AM.
Figure 35: Interpolation curves of test accuracy for modified decision lists with AM.
Figure 36: Interpolation curves of train accuracy for modified decision lists with WM.
Figure 37: Interpolation curves of test accuracy for modified decision lists with WM.
Figure 38: Interpolation curves of train accuracy for modified decision lists with AM by use of split dataset.
Figure 39: Interpolation curves of test accuracy for modified decision lists with AM by use of split dataset.
Figure 40: Interpolation curves of train accuracy for modified decision lists with WM by use of split dataset.
Figure 41: Interpolation curves of test accuracy for modified decision lists with WM by use of split dataset.
Figure 42: Interpolation curves of test accuracy with WM for MNIST [35] dataset.
Figure 43: Interpolation curves of test accuracy with AM for MNIST [35] dataset.